OpenStudy (anonymous):

The only solution to x+1=72/x is x = -9. TRUE OR FALSE

OpenStudy (anonymous):

true

OpenStudy (hba):

it will have two answers so false

OpenStudy (anonymous):

true :)

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Im a little confused everyone..is it true or false??

OpenStudy (anonymous):

True

OpenStudy (hba):

the other answer is 8

OpenStudy (anonymous):

x+1=72/x => x^2 + x - 72 = 0 (x-8)(x+9) = 0 => x =8, x=-9

OpenStudy (anonymous):

It may have two solution but the solution x=-9 is true

OpenStudy (anonymous):

im gonna go with true :)

OpenStudy (anonymous):

its false...definitely false...because its not the "ONLY" solution

OpenStudy (anonymous):

agree @tapan